OpenStudy (aleah54):

FAN AND MEDAL A barrel of tomato sauce has spilled on a tile floor. The sauce flow can be expressed with the function r(t) = 2t, where t represents time in minutes and r represents how far the sauce is spreading. The spilled sauce is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr2. Part A: Find the area of the circle of spilled sauce as a function of time, or A[r(t)]. Show your work. (6 points) Part B: How large is the area of spilled sauce after 5 minutes? You may use 3.14 to approximate π in this problem. (4 points)

Vocaloid (vocaloid):

for part A, we want to rewrite r in terms of t

Vocaloid (vocaloid):

r = 2t then we substitute that into the area equation A = pi * r^2 = pi * (2t)^2

Vocaloid (vocaloid):

for part B, you just let t = 5 and solve for A

Vocaloid (vocaloid):

let me know if you have additional questions

OpenStudy (aleah54):

I still don't get part B @vocaloid

Vocaloid (vocaloid):

t = 5 A = pi * (2t)^2 = ?

OpenStudy (aleah54):

so do i put 4t? @vocaloid

Vocaloid (vocaloid):

no

Vocaloid (vocaloid):

if t = 5 what does pi * (2t)^2 equal?

OpenStudy (aleah54):

seven @vocaloid

Vocaloid (vocaloid):

no

Vocaloid (vocaloid):

what is (2t)^2?

OpenStudy (aleah54):

9 i think

Vocaloid (vocaloid):

where are you getting these numbers??

Vocaloid (vocaloid):

if t = 5 what is 2t?

OpenStudy (aleah54):

5 times 2

Vocaloid (vocaloid):

which equals...?

OpenStudy (aleah54):

oh i multiplied wrong it's 10

Vocaloid (vocaloid):

good now, what is 10^2?

OpenStudy (aleah54):

the ^ sign means to divide right

Vocaloid (vocaloid):

no, it's an exponent

Vocaloid (vocaloid):

10^2 = 10 * 10 = ?

OpenStudy (aleah54):

oh so it's 20

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

what is 10 * 10?

OpenStudy (aleah54):

100

Vocaloid (vocaloid):

good now what is 100 * 3.14?

OpenStudy (aleah54):

314

Vocaloid (vocaloid):

good, that's your final answer

OpenStudy (aleah54):

thank you so much